# The resolution quiver of Nakayama algebras which are minimal Auslander-Gorenstein

### v1

@@ Abstract.
Let $A$ be a Nakayama algebra. Using Ringel’s resolution quiver, we give a criterion to decide whether $A$ is minimal Auslander-Gorenstein. The criterion strongly relies on the parity of the selfinjective dimension of $A$.
@@ 1. Introduction
Following Iyama [4], an algebra is said to be higher Auslander if its global dimension is finite and bounded by its dominant dimension. Higher Auslander algebras arise in cluster-tilting theory. Following Iyama and Solberg [5], an algebra is said to be minimal Auslander-Gorenstein provided that its global dimension is finite and bounded by its dominant dimension. Minimal Auslander-Gorenstein algebras are a special class of Auslander-Gorenstein algebras.
Nakayama algebras are one of most basic classes of algebras in representation theory of algebras. An algebra is called a Nakayama algebra if it is a homomorphic image of the path algebra of an oriented cycle. A connected Nakayama algebras is either linear or else cyclic.
@@
The syzygy filtered algebras for cyclic Nakayama algebras is introduced by Sen [12]. It plays a role in constructing cyclic Nakayama algebras which are higher Auslander in [13]. The reversed syzygy filtered algebras for all Nakayama algebras is introduced in [15]. It provides a systematic construction of cyclic Nakayama algebras which are minimal Auslander-Gorenstein and higher Auslander.
Inspired by their work, we study syzygy filtered algebras for all Nakayama algebras. Then we give a resolution quiver criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein. This criterion strongly depends on the parity of the selfinjective dimension.
Let $A$ be a Nakayama algebra with translation $\tau$. For an $A$-module $M$, denote by $P(M)$ the projective cover and by $I(M)$ the injective envelope of $M$. The resolution quiver $R(A)$ of $A$ is defined as follows. The vertex set of $R(A)$ is the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to $\tau\operatorname{soc}P(S)$ for each simple $A$-module $S$. A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise. This assigns colors to the vertices of resolution quivers.
The weight of a cycle $C$ in $R(A)$ is the sum of $c_{a}/n$, where $a$ runs through all vertices in $C$ and $n$ is rank of $A$. The weight is an integer and all cycles in $R(A)$ have the same weight. The weight of any cycle in $R(A)$ is called the weight of $R(A)$.
A simple $A$-module $S$ said to be cyclic if $S$ lies on cycle in $R(A)$. For any simple $A$-module $S$, denote by $\operatorname{dist}S$ the minimal integer $d$ such that the $d$-th successor of $S$ is cyclic in $R(A)$. Let $\operatorname{dist}A$ be the maximum of the numbers $\operatorname{dist}S$, where $S$ runs through all simple $A$-modules.
@@
Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of odd selfinjective dimension if and only if
@@
    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
@@
Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of even selfinjective dimension if and only if
@@
    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
@@
    every cyclic vertex has at most $2$ predecessor in $R(A)$;
@@
Throughout this paper, $k$ is a fixed algebraically closed field. All algebras are finite dimensional associative algebras over $k$. All modules are finite dimensional left modules unless otherwise stated.
The paper is organized as follows. In Section 2, we recall some basic properties of Nakayama algebras. In Section 3, we study syzygy filtered algebra for all Nakayama algebras. Then we prove the criterion in Section 4. Some applications and examples is given in Section 5.
@@ Acknowledgements
The author is grateful to Professor Shijie Zhu for his comments and suggestions. This work is supported by the National Natural Science Foundation of China (No. 1257011205).
@@ 2. Preliminary
We refer to [1] for quivers and their representations.
Given an integer $n$, we denote by $\Delta_{n}$ the quiver on $n$ vertices containing a single directed cycle through all vertices; see Figure 1. Denote by $k\Delta_{n}$ the path algebra of the quiver $\Delta_{n}$.
@@
Let $A$ be a Nakayama algebra of rank $n$. For each vertex $a$ in $\Delta_{n}$, denote by $S(a)$ the simple $A$-module, $P(a)$ the projective $A$-module and $I(a)$ the injective $A$-module at $a$. Set $\tau S(a)=S(b)$ if there is an arrow from $a$ to $b$ in $\Delta_{n}$. The map $\tau$ yields a translation of simple $A$-modules. If $A$ is a cyclic Nakayama algebra, then $\tau$ coincides with the Auslander-Reiten translation.
@@
$c(A)=(c_{a},c_{\tau a},\cdots,c_{\tau^{n-1}a})$
is called a Kupisch series of $A$. Two Nakayama algebras $A$ and $B$ are isomorphic if and only if $c(A)$ and $c(B)$ are equal up to a power of cyclic permutation.
Denote by $R(A)$ the resolution quiver of $A$. The vertices of $R(A)$ are the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to
@@
for each simple $A$-module $S$; see also [2]. A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise. This assigns colors to the vertices of $R(A)$.
@@
An algebra $A$ is said to be Gorenstein if the left selfinjective dimension and right selfinjective dimension of $A$ are finite. For any Gorenstein algebra, two selfinjective dimensions coincide; see [17]. It is known that a Nakayama algebra is Gorenstein if and only if its left selfinjective dimension is finite.
We need the following; see [16, 1.2] and [12, 4.13].
@@
    AA is Gorenstein if and only if all cycles in $R(A)$ are black;
@@
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. A vertex $S$ in $R(A)$ is said to be a leaf if there is no arrow ending with $S$.
We need the following; see [6, 3.1].
@@
Let $S$ be a simple $A$-module. Then $\mathrm{id}_{A}\,S=1$ if and only if $S$ is a leaf in $R(A)$.
@@ 3. Syzygy filtered algebra
In the present section, we study the syzygy filtered algebra for all Nakayama algebras and give some necessary results for the proof of the main result.
Let $A$ be a Nakayama algebra of rank $n$ and let $R(A)$ be the resolution quiver of $A$. Following [11, A.6] the map $\gamma$ induces a one-to-one correspondence between the set of all black vertices and the set of all non-leaf vertices in $R(A)$. For any non-leaf vertex $S$ in $R(A)$, let $T$ be the unique black vertex such that $\gamma(T)=S$. Consider the exact sequence
$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0.$ (3.1)
The base module $\Delta(S)$ is the quotient module of $P(S)$ of minimal length such that the translation of its socle is a non-leaf vertex.
@@
form a semibrick in the category $\operatorname{mod}A$ of all finite dimensional $A$-modules. The filtration category $\mathcal{F}$ of $\mathcal{B}$ is an exact Abelian subcategory of $\operatorname{mod}A$. The category $\mathcal{F}$ contains the projective cover $\mathcal{P}$ of all base modules and the second syzygy of all finite dimensional $A$-modules.
The syzygy filtered algebra $\varepsilon(A)$ is the endomorphism algebra of $\mathcal{P}$, that is,
@@
It is a Nakayama algebra of rank at most $n$. The equality holds if and only if $A$ is selfinjective. The category $\operatorname{mod}\varepsilon(A)$ is equivalent to $\mathcal{F}$. The algebra $\varepsilon(A)$ has finite global dimension if and only if $A$ has finite global dimension.
For a Nakayama algebra $A$, denote by $\chi(A)$ the number of isoclasses of simple projective $A$-modules. By (3.2), we have the following.
@@
Let $A$ be a Nakayama algebra. Then $\chi(\varepsilon{A})-\chi(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2$.
For any integer $i\geq 0$, denote by $\varepsilon^{i}(A)$ the $i$-th iterated syzygy filtered algebra of $A$. We have the following.
Corollary 3.2.
For any $i\geq 1$, $\chi(\varepsilon^{i}{A})-\chi(\varepsilon^{i-1}A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i$.
There is an integer $r$ such that $\varepsilon^{r}(A)$ is a selfinjective Nakayama algebra. Then $\chi(\varepsilon^{r}{A})$ is equal to the number of isoclasses of simple $A$-modules of even projective dimension. This gives an alternative proof of Madsen’s criterion [6], which states that a Nakayama algebra has finite global dimension if and only if it has a simple module of even projective dimension.
The projective objects in $\mathcal{F}$ are precisely the projective $A$-modules which are contained in $\mathcal{F}$. However, the injective objects in $\mathcal{F}$ are more complicated.
Following [7], the defect $\operatorname{def}A$ of an algebra $A$ is the number of isoclasses of indecomposable injective non-projective $A$-modules. If $A$ is a Nakayama algebra, then $\operatorname{def}A$ is equal to the number of red vertices in its resolution quiver. By (3.2) we have following.
@@
For any $i\geq 0$, $\operatorname{def}\varepsilon^{i}(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i+1$.
Let $S$ be a non-leaf vertex in $R(A)$. By [11, A.6] there is a unique vertex $T$ such that $P(T)$ is injective and $\gamma(T)=S$. Consider the exact sequence
$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\overset{f}{\to}I(\tau^{-1}T)\to 0.$ (3.3)
We have the following; compare [12, 3.18].
@@
If $T$ is a non-leaf vertex, the module $I(\tau^{-1}T)$ is projective. The sequence (3.3) splits and $J(\tau^{-1}S)$ is an indecomposable injective $A$-module contained in $\mathcal{F}$.
If $T$ is a leaf, then $I(\tau^{-1}T)$ is not projective and $\operatorname{Ker}f$ is nonzero. Since $P(T)$ is an injective $A$-module, $J(\tau^{-1}S)$ is nonzero. Then $J(\tau^{-1}S)$ is an indecomposable object in $\mathcal{F}$. Since $\operatorname{Ker}f$ is a proper submodule of a base module and $P(T)$ is an injective $A$-module, we know that $J(\tau^{-1}S)$ is an injective object in $\mathcal{F}$.
The number of isoclasses of injective objects in $\mathcal{F}$ is equal to the number of non-leaf vertices in $R(A)$. If $S$ and $S^{\prime}$ are not isomorphic, $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic. Then the first conclusion holds.
Note that $J(\tau^{-1}S)$ is a non-projective object in $\mathcal{F}$ if and only if $\Delta(S)$ is a leaf in $R(\varepsilon{A})$. This happens if and only if the injective dimension of $S$ is $3$. ∎
The translation $\tau$ induces a translation $\tau_{\varepsilon}$ of simple objects in $\mathcal{F}$. For any vertex $\Delta(S)$ in $R(\varepsilon{A})$, by [11, B.4] we have
@@
The assignment $\Delta(S)\mapsto S$ yields an injective map
$\varphi\colon R(\varepsilon{A})\to R(A)$
of resolution quivers. Denote by $R(A)^{\prime}$ the image of $\varphi$. We mention that $\varphi$ preserves the weight of any cycle. However, $\varphi$ may not preserve the vertex colors.
A vertex $T$ in $R(A)$ is said to be *-black if $\tau^{-1}T$ is black, and $T$ is said to be *-red if $\tau^{-1}T$ is red. For each leaf $S$ in $R(A)^{\prime}$, there is a unique black leaf $T_{0}$ in $R(A)$ such that $\gamma(T_{0})=S$ and a unique *-black leaf $T_{1}$ in $R(A)$ such that $\gamma(T_{1})=S$. For any resolution quiver $R$, denote by $\mathrm{Leaf}\mbox{-}R,\mathrm{BLeaf}\mbox{-}R$ and $\mathrm{*BLeaf}\mbox{-}R$ the set of all leaf vertices, black leaf vertices and *-black leaf vertices, respectively. There is an injective map
$\varphi_{0}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{BLeaf}\mbox{-}R(A)$
sending $S$ to $T_{0}$ and an injective map
$\varphi_{1}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{*BLeaf}\mbox{-}R(A)$
sending $S$ to $T_{1}$. Consider the following two conditions.
@@
    For each black leaf in $R(A)$, each predecessor of its successor is a leaf.
@@
    For each *-black leaf in $R(A)$, each predecessor of its successor is a leaf.
@@
    $\varphi$ preserves vertex colors;
@@
    The number of red vertices in $R(\varepsilon{A})$ and that $R(A)^{\prime}$ are equal.
@@
$|\mathrm{Red}\mbox{-}R(A)^{\prime}|=|\mathrm{Red}\mbox{-}R(A)|-|\mathrm{RLeaf}\mbox{-}R(A)|=|\mathrm{BLeaf}\mbox{-}R(A)|$
@@
$|\mathrm{Red}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(A)^{\prime}|\leq|\mathrm{BLeaf}\mbox{-}R(A)|.$
@@
$(2)\iff(3)$ Let $P$ be an indecomposable projective object in $\mathcal{F}$. If $P$ is minimal in $\operatorname{mod}A$, so is $P$ in $\mathcal{F}$. This shows that $\varphi$ preserves red vertices. Then $\varphi$ preserves vertex colors if and only if the numbers of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal. ∎
@@
Let $A$ be a Nakayama algebra satisfying (B). Then $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Moreover, we have
@@
Since $A$ satisfies (B), then $\varphi$ preserves vertex colors. Recall that $A$ has finite global dimension if and only if $\varepsilon(A)$ has finite global dimension. Then by Lemma 2.4 the algebra $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Note that the selfinjective dimension of $A$ is not $1$. The equation (3.4) follows from [12, 3.23]. ∎
@@
Similarly, we have the following.
Proposition 3.9.
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    AA satisfies (B);
  2. (2)
    $\varphi$ preserves vertex *-colors;
  3. (3)
    The number of *-red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal.
@@
Note that $J(4)$ is isomorphic to $P(1)$ and $J(1)$ is an indecomposable injective non-projective object in $\mathcal{F}$.

### v2

@@ Abstract.
Let $A$ be a Nakayama algebra. Using Ringel’s resolution quiver, we give a criterion to determine whether $A$ is a minimal Auslander-Gorenstein algebra. The criterion strongly relies on the parity of the selfinjective dimension of $A$.
@@ 1. Introduction
Following Iyama [4], an algebra is said to be higher Auslander provided that its global dimension is finite and bounded by its dominant dimension. Higher Auslander algebras generalize the classical Auslander algebras. Following Iyama and Solberg [5], an algebra is said to be minimal Auslander-Gorenstein provided that its selfinjective dimension is finite and bounded by its dominant dimension. Minimal Auslander-Gorenstein algebras are the Gorenstein analogues of higher Auslander algebras. They are a special class of Auslander-Gorenstein algebras.
Nakayama algebras are one of the most fundamental classes of algebras in the representation theory of algebras. An algebra is called a Nakayama algebra if it is a homomorphic image of the path algebra of an oriented cycle. The indecomposable modules over Nakayama algebras are uniserial. A connected Nakayama algebra is either linear or cyclic.
@@
The syzygy filtered algebra for cyclic Nakayama algebras is introduced by Sen [13]. It plays a role in constructing cyclic Nakayama algebras which are higher Auslander algebras [14]. The reversed syzygy filtered algebra for all Nakayama algebras is introduced in [16]. It provides a systematic method to construct cyclic Nakayama algebras which are higher Auslander and minimal Auslander-Gorenstein. A classification of defect one cyclic Nakayama algebras is given in [7], and a classification of concave linear Nakayama algebras which are higher Auslander is given in [12].
Inspired by their work, we study the syzygy filtered algebra for all Nakayama algebras. Then we give a resolution quiver criterion to determine whether a Nakayama algebra is a minimal Auslander-Gorenstein algebra. This criterion strongly depends on the parity of the selfinjective dimension.
Let $A$ be a Nakayama algebra with translation $\tau$ on the simple $A$-modules. For an $A$-module $M$, denote by $\ell(M)$ the composition length, by $P(M)$ the projective cover, and by $I(M)$ the injective envelope of $M$. The resolution quiver $R(A)$ of $A$ is defined as follows. The vertices of $R(A)$ are the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to $\tau\operatorname{soc}P(S)$ for each simple $A$-module $S$. A simple $A$-module $S$ is said to be red if the projective dimension of $S$ is equal to $1$, and black otherwise. This assigns a color to each vertex of the resolution quiver.
A simple $A$-module $S$ is said to be cyclic if $S$ lies on a cycle in $R(A)$, and a leaf if there is no arrow in $R(A)$ ending at $S$. For any simple $A$-module $S$, we denote by $\operatorname{dist}S$ the minimal integer $d$ such that the $d$-th successor of $S$ is cyclic. Let $\operatorname{dist}A$ be the maximum of $\operatorname{dist}S$ over all simple $A$-modules $S$.
The weight of $R(A)$ is the sum of $\ell(P(S))/nc$, where $n$ is the rank of $A$, $c$ is the number of cycles in $R(A)$, and $S$ runs through all cyclic vertices in $R(A)$. It turns out that a Nakayama algebra has finite global dimension if and only if its resolution quiver is connected of weight $1$.
@@
Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of odd selfinjective dimension if and only if
@@
    $\operatorname{dist}S=\operatorname{dist}A>0$ for every leaf $S$ in $R(A)$;
@@
Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of even selfinjective dimension if and only if
@@
    $\operatorname{dist}S=\operatorname{dist}A>0$ for every leaf $S$ in $R(A)$;
@@
    every cyclic vertex has at most $2$ predecessors in $R(A)$;
@@
Throughout this paper, $k$ is a fixed algebraically closed field. All algebras are finite dimensional associative algebras over $k$. All modules are finite dimensional left modules unless otherwise stated. We refer to [1] for basics on finite dimensional algebras and their representations.
The paper is organized as follows. In Section 2, we recall some basic properties of Nakayama algebras and their resolution quivers. In Section 3, we investigate the syzygy filtered algebra for all Nakayama algebras. Then we prove the main result in Section 4. Some applications and examples are given in Section 5.
@@ Acknowledgements
The author is grateful to Professor Shijie Zhu for his comments and suggestions. This work is supported by the National Natural Science Foundation of China (No. 12571036 and 11801141).
@@ 2. Preliminary
Given an integer $n$, we denote by $\Delta_{n}$ the quiver on $n$ vertices containing a single directed cycle through all vertices; see Figure 1. Denote by $k\Delta_{n}$ the path algebra of $\Delta_{n}$.
@@
Let $A$ be a Nakayama algebra of rank $n$. For each vertex $a$ in $\Delta_{n}$, denote by $S(a)$ the simple $A$-module, by $P(a)$ the projective $A$-module, and by $I(a)$ the injective $A$-module at $a$. Set $\tau(a)=b$ if there is an arrow from $a$ to $b$ in $\Delta_{n}$. This induces a translation $\tau$ on the simple $A$-modules. If $A$ is a cyclic Nakayama algebra, then $\tau$ coincides with the Auslander-Reiten translation [1].
@@
$c(A)=(c_{1},c_{2},\cdots,c_{n})$
is called a Kupisch series of $A$. Two Nakayama algebras are isomorphic if and only if their Kupisch series are equal up to cyclic permutation.
Following [10], the resolution quiver $R(A)$ of $A$ is the successor graph of the function
@@
on the set of all isoclasses of simple $A$-modules; see also [2]. A simple $A$-module $S$ is said to be red if the projective dimension of $S$ is equal to $1$, and black otherwise. A projective $A$-module $P$ is said to be minimal if its top module is a direct sum of black simple $A$-modules.
@@
An algebra $A$ is said to be Gorenstein if its left selfinjective dimension and right selfinjective dimension are finite. If $A$ is a Gorenstein algebra, then its left and right selfinjective dimensions are equal; see [18]. It is known that a Nakayama algebra is Gorenstein if and only if its left selfinjective dimension is finite.
We need the following; see [17, 1.2] and [13, 4.13].
@@
    AA is Gorenstein if and only if all cyclic vertices in $R(A)$ are black;
@@
Let $A$ be a Nakayama algebra. A simple $A$-module $S$ is said to be a leaf if there is no arrow in $R(A)$ ending at $S$. We need the following; see [6, 3.1].
@@
Let $S$ be a simple $A$-module. Then
  1. (1)
    $I(S)$ is projective if and only if $\tau S$ is a non-leaf in $R(A)$;
  2. (2)
    $\mathrm{id}_{A}\,S=1$ if and only if $S$ is a leaf in $R(A)$.
@@ 3. Syzygy filtered algebra
In this section, we study the syzygy filtered algebra for all Nakayama algebras. This generalizes Sen’s construction for cyclic Nakayama algebras.
Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. By [11, A.6] the map $\gamma$ induces a one-to-one correspondence between the set of all black vertices and the set of all non-leaf vertices in $R(A)$. For any non-leaf vertex $S$ in $R(A)$, let $T$ be the unique black vertex such that $\gamma(T)=S$. Consider the exact sequence
$0\to\Delta(S)\to P(\tau T)\overset{\alpha}{\to}P(T)\to T\to 0,$ (3.1)
where $\alpha\in A$ is the arrow starting at $T$. The kernel $\Delta(S)$ is called a base module. It is the quotient module of $P(S)$ of minimal length such that the translation of its socle is a non-leaf vertex.
@@
form a semibrick in the category $\operatorname{mod}A$ of all finite dimensional $A$-modules. The filtration category $\mathcal{F}$ of $\mathcal{B}$ is an exact Abelian subcategory of $\operatorname{mod}A$. The category $\mathcal{F}$ contains the projective cover of every base module and the second syzygy of every finite dimensional $A$-module.
Denote by $\mathcal{P}$ the projective covers of all base modules. The syzygy filtered algebra $\varepsilon(A)$ of $A$ is the endomorphism algebra of $\mathcal{P}$, that is,
@@
It is a Nakayama algebra of rank less than or equal to the rank of $A$. The equality holds if and only if $A$ is selfinjective. The category $\operatorname{mod}\varepsilon(A)$ of all finite dimensional $\varepsilon(A)$-modules is equivalent to $\mathcal{F}$. The syzygy filtered algebra $\varepsilon(A)$ has finite global dimension if and only if $A$ has finite global dimension.
Given a Nakayama algebra $A$, denote by $\chi(A)$ the number of isoclasses of simple projective $A$-modules. Note that $\chi(A)$ equals the number of occurrences of $1$ in the Kupisch series of $A$. For any integer $i\geq 0$, denote by $\varepsilon^{i}(A)$ the $i$-th iterated syzygy filtered algebra of $A$. By (3.2), we have the following.
@@
Let $A$ be a Nakayama algebra and $i\geq 1$. Then $\chi(\varepsilon^{i}{A})-\chi(\varepsilon^{i-1}A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i$.
For any $i\geq 1$, the rank of $\varepsilon^{i}(A)$ is no more than the rank of $\varepsilon^{i-1}(A)$. There is an integer $m$ such that $\varepsilon^{m}(A)$ is selfinjective. Then $\chi(\varepsilon^{m}{A})$ is equal to the number of isoclasses of simple $A$-modules of even projective dimension. This gives an alternative proof of Madsen’s criterion [6], which states that a Nakayama algebra $A$ has finite global dimension if and only if there is a simple $A$-module of even projective dimension.
Following [7], the defect $\operatorname{def}A$ of an algebra $A$ is the number of isoclasses of indecomposable injective non-projective $A$-modules. If $A$ is a Nakayama algebra, the defect of $A$ is equal to the number of red vertices in the resolution quiver of $A$. By (3.2) we have following.
@@
Let $A$ be a Nakayama algebra and $i\geq 0$. Then $\operatorname{def}\varepsilon^{i}(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i+1$.
The projective objects in $\mathcal{F}$ are precisely the projective $A$-modules contained in $\mathcal{F}$. The injective objects in $\mathcal{F}$ are more complicated. Let us give a description.
Let $S$ be a non-leaf vertex in $R(A)$. By [11, A.6] there is a unique vertex $T$ such that $\tau^{-1}T$ is black and $\gamma(T)=S$. Consider the exact sequence
$0\to J(\tau^{-1}S)\to P(T)\overset{w}{\to}P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0,$ (3.3)
where $w\in A$ is the concatenation of the maximal nonzero path ending at $\tau^{-1}T$ and the arrow starting at $\tau^{-1}T$.
We have the following; compare [13, 3.18].
@@
If $T$ is a non-leaf vertex, then $I(\tau^{-1}T)$ is projective and $w$ is zero. It follows that $J(\tau^{-1}S)$ is an indecomposable injective $A$-module contained in $\mathcal{F}$.
If $T$ is a leaf, then $I(\tau^{-1}T)$ is not projective and $w$ is nonzero. Since $P(T)$ is an injective $A$-module, $J(\tau^{-1}S)$ is an indecomposable object in $\mathcal{F}$. Since the image of $w$ is a proper submodule of a base module, we infer that $J(\tau^{-1}S)$ is an injective object in $\mathcal{F}$.
If $S$ and $S^{\prime}$ are not isomorphic, the socle modules of $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic. Then $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic. Note that the number of isoclasses of indecomposable injective objects in $\mathcal{F}$ is equal to the number of non-leaf vertices in $R(A)$. Then the assignment is a bijection.
By Lemma 2.5 $J(\tau^{-1}S)$ is a non-projective object in $\mathcal{F}$ if and only if $\Delta(S)$ is a leaf in $R(\varepsilon{A})$. This happens if and only if the injective dimension of $S$ is $3$. ∎
The translation $\tau$ induces a translation $\tau_{\varepsilon}$ one the simple objects in $\mathcal{F}$. For any vertex $\Delta(S)$ in $R(\varepsilon{A})$, by [11, B.4] we have
@@
The assignment $\Delta(S)\mapsto S$ yields an injective map of resolution quivers
$\varphi\colon R(\varepsilon{A})\to R(A).$
We denote by $R(A)^{\prime}$ the image of $\varphi$. Then $R(A)^{\prime}$ is obtained from $R(A)$ by removing all leaf vertices. Note that $\varphi$ preserves the weight. However, $\varphi$ may not preserve the vertex colors.
A simple $A$-module $S$ is called *-black if $\tau^{-1}S$ is black, and *-red otherwise. For any resolution quiver $R$, denote by $\mathrm{Leaf}\mbox{-}R,\mathrm{BLeaf}\mbox{-}R$, and $\mathrm{*BLeaf}\mbox{-}R$ the set of all leaf vertices, black leaf vertices, and *-black leaf vertices, respectively.
For each leaf $S$ in $R(A)^{\prime}$, there is a unique black leaf $T_{0}$ in $R(A)$ such that $\gamma(T_{0})=S$, and a unique *-black leaf $T_{1}$ in $R(A)$ such that $\gamma(T_{1})=S$. There is an injective map
$\varphi_{0}\colon\mathrm{Leaf}\mbox{-}R(\varepsilon{A})\to\mathrm{BLeaf}\mbox{-}R(A)$
sending $\Delta(S)$ to $T_{0}$ and an injective map
$\varphi_{1}\colon\mathrm{Leaf}\mbox{-}R(\varepsilon{A})\to\mathrm{*BLeaf}\mbox{-}R(A)$
sending $\Delta(S)$ to $T_{1}$. Consider the following two conditions.
@@
    For each black leaf in $R(A)$, its successor is a leaf in $R(A)^{\prime}$.
@@
    For each *-black leaf in $R(A)$, its successor is a leaf in $R(A)^{\prime}$.
@@
    $\varphi$ preserves the vertex colors;
@@
    the number of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal;
  4. (4)
    any minimal projective object in $\mathcal{F}$ is a minimal projective $A$-module.
@@
$|\mathrm{Red}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(\varepsilon{A})|\leq|\mathrm{BLeaf}\mbox{-}R(A)|$
@@
$|\mathrm{BLeaf}\mbox{-}R(A)|=|\mathrm{Leaf}\mbox{-}R(A)|-|\mathrm{RLeaf}\mbox{-}R(A)=|\mathrm{Red}\mbox{-}R(A)^{\prime}|.$
@@
$(2)\iff(3)$ Let $P$ be an indecomposable projective object in $\mathcal{F}$. If $P$ is minimal projective in $\operatorname{mod}A$, then $P$ is minimal projective in $\mathcal{F}$. This shows that $\varphi$ preserves red vertices. Then $\varphi$ preserves the vertex colors if and only if the number of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal.
$(2)\iff(4)$ This is obvious. ∎
Similarly, we have the following.
Proposition 3.5.
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    AA satisfies (B);
  2. (2)
    $\varphi$ preserves the vertex *-colors;
  3. (3)
    the number of *-red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal;
  4. (4)
    any injective projective object in $\mathcal{F}$ is an injective projective $A$-module.
@@
Let $A$ be a Nakayama algebra satisfying (B). Then $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Furthermore, we have
@@
Since $A$ satisfies (B), then $\varphi$ preserves the vertex colors. Recall that $A$ has finite global dimension if and only if $\varepsilon(A)$ has finite global dimension. Then by Lemma 2.4 the algebra $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein.
If $\mathrm{id}_{A}\,A=1$, then $A$ is hereditary by Lemma 2.4. There is a black leaf in $R(A)$. This is impossible. The equation (3.4) follows from [13, 3.23]. ∎
@@
Note that $J(4)$ is an injective projective $A$-module, and $J(1)$ is a non-projective $A$-module.